sum of two numbers is 49 and the difference is 25
Answer by arunpaul(104)

Let the number be x and y
as per the question
x+y = 49 - (eq1)
x-y = 25 - (eq2)
adding eq1 and eq2 we have
x+y = 49
x-y = 25
------------
2x = 74
x = 74/2 = 37
putting the value in equation 1 we have
37 + y = 49
y = 49-37 = 12
so x = 37 and y = 12 (Ans)